OpenStudy (anonymous):

[WILL FAN ANYONE THAT HELPS] ][8th Grade Mathematics C] Simplify the following expression. 9[squared by 6] ÷ 9[squared by negative 3] A 9[squared by negative 9] B 9[squared by 9] C 9[squared by negative 18] D9[squared by 3]

OpenStudy (triciaal):

can you rewrite the expression in the question ?

OpenStudy (anonymous):

Subtract the exponent of the denominator from the exponent of the numerator\[6-(-3)=6+3=9\]

OpenStudy (anonymous):

Raise the base number (9) to the value you got from subtracting the exponent of the denominator from the exponent of the numerator (9) to get your answer:\[9^9\]

OpenStudy (anonymous):

That's choice B

OpenStudy (triciaal):

@samarasamara can you rewrite your interpretation please

OpenStudy (triciaal):